# Large Language Model Supply Chain: A Research Agenda

### v2

@@ 5.4. Model: Deploy & Maintenance
Challenge I [SE Perspective]: Model Drift. The challenge of identifying and quantifying model drift in LLMs is considerable (Madaan et al., 2023; CITTERIO, 2022; Oroy and Evan, 2024; Li et al., 2024c). LLMs are trained on vast datasets that are supposed to represent the linguistic diversity of their intended application domain. However, as the language evolves or the model is applied to slightly different contexts, the ability to remain relevant and consistent can shift in subtle ways. Recent research (Madaan et al., 2023; CITTERIO, 2022; Broscheit et al., 2022) emphasizes the need for sophisticated tools that can detect not only overt drifts in language usage but also more nuanced shifts in sentiment, context, or cultural references. These tools must be capable of parsing the complexities of human language, requiring ongoing refinement and adaptation to new linguistic phenomena.
  • Opportunity: Model Drift Monitoring. The realm of drift monitoring in LLMs presents a fertile ground for innovation and development. There is a significant opportunity to create and refine tools that can accurately detect and measure drift in various dimensions, from language usage to sentiment and contextual nuances. Furthermore, integrating these drift monitoring tools into the model development and deployment lifecycle can provide ongoing insights into model performance (CITTERIO, 2022; Madaan et al., 2023), enabling timely adjustments and enhancements. This proactive approach to managing model drift not only ensures the sustained relevance and accuracy of LLMs but also opens new avenues for research in understanding and mitigating the subtleties of language evolution in artificial intelligence.
@@ 6.1. Application: LLM Chatbots and Applications
Current State. Platforms like the GPT Store (OpenAI, 2024) are emerging as centralized hubs where developers can publish their LLM-powered apps (i.e., GPTs), and users can discover and access these tools to fulfill a wide array of tasks and objectives. This ecosystem draws parallels with mobile app stores (Fu et al., 2013; Lee and Raghu, 2014; Martin et al., 2016), aiming to create a curated, secure, and user-friendly environment for LLM-driven apps. The proliferation of LLM chatbots and apps is catalyzing innovation by lowering the barriers to entry for developers and providing them with platforms to build and deploy LLM-powered solutions. However, this rapid growth and accessibility also introduce new vulnerabilities and governance challenges. As these platforms evolve, they must grapple with novel security threats specific to LLM applications (Hou et al., 2024a; Su et al., 2024), such as prompt injection attacks and the potential misuse of advanced features like function calling. Furthermore, the unique nature of LLM apps, which can generate and manipulate content in real time, presents unprecedented challenges in terms of quality control, ethical considerations, and regulatory compliance.
Challenge I [S&P Perspective]: Security Vulnerabilities in LLM Apps. LLM-powered apps face significant security challenges, particularly in the realms of prompt injection attacks, misuse of function calling features, and prompt leakage (Wu et al., 2024a; OWASP, 2024b). Prompt injection attacks can manipulate the behavior of LLM chatbots, potentially leading to unauthorized actions or the disclosure of sensitive information (Liu et al., 2023a; Greshake et al., 2023; Liu et al., 2023e). Function calling features, while powerful, introduce new attack vectors. Attackers may exploit these features to extract parameters and sensitive information about the underlying tools, hijack data flows, or conductresource exhaustion attacks (OWASP, 2024b) by overloading the system with malicious function calls (OWASP, 2024b). Additionally, prompt leakage poses a significant risk, where sensitive information embedded in prompts may be inadvertently revealed through the model’s responses, potentially compromising user privacy and system security (Hou et al., 2024a; Su et al., 2024).
@@ 6.2. Application: Autonomous LLM-based Agents
Current State. The current landscape of ALAs is characterized by rapid advancement and diverse applications across multiple domains. Recent developments have seen the emergence of more sophisticated agent architectures that combine LLMs with external tools, knowledge bases, and decision-making frameworks (Guo et al., 2024; Xi et al., 2023). These agents are increasingly capable of performing complex, multi-step tasks with minimal human intervention. For instance, in the field of software development, ALAs are being used for code generation, debugging, and even system design (Jin et al., 2024; Liu et al., 2024b; Huang et al., 2024). However, alongside these advancements, there is growing concern about the ethical implications and potential risks associated with increasingly autonomous agents (He et al., 2024c; Li et al., 2024f; Zhang et al., 2024b). Current research is focused on enhancing the reliability, transparency, and controllability of ALAs, with particular emphasis on developing robust safety measures and alignment techniques to ensure agent behaviors remain consistent with human intentions and values.
Challenge I [SE Perspective]: Agent Optimization. From a software engineering perspective, a significant challenge in ALA development is the optimization of agent performance through efficient workflow design and orchestration (Sun et al., 2024b). This challenge falls under the emerging field of Software Engineering for LLM (SE4LLM). The complexity of tasks that ALAs can handle requires careful consideration of how different components interact, how information flows between stages of processing, and how to balance autonomy with controlled behavior. Effective agent optimization involves designing modular and reusable components, implementing efficient error handling and recovery mechanisms, and creating adaptive workflows that can adjust to varying task requirements and environmental conditions (Liu et al., 2024b; Huang et al., 2024).
  • Opportunity: Workflow Orchestration Frameworks. This challenge presents an opportunity to develop sophisticated workflow orchestration frameworks specifically tailored for ALAs. Such frameworks could provide tools for visual workflow design, automated performance analysis, and dynamic workflow adjustment based on real-time feedback. Research in this area could focus on creating best practices for agent workflow design (Liu et al., 2024b), developing metrics for agent efficiency and effectiveness, and exploring ways to optimize the interaction between LLM components and external tools or data sources.
@@ 7. Conclusion
This paper presents a comprehensive research agenda for the LLM supply chain, identifying the complex challenges and opportunities arising from the rapid advancement of Large Language Models. By examining the LLM supply chain through the dual lenses of Software Engineering and Security & Privacy, we have provided a structured approach to understanding and advancing this critical field. Our analysis spans the infrastructure layer, foundation model layer, and downstream application ecosystem, presenting visions for robust and secure LLM development, reviewing current practices, and identifying key challenges and research opportunities. This work bridges the existing research gap in systematically understanding the multifaceted issues within the LLM supply chain. The proposed research agenda offers valuable insights to guide future efforts, fostering innovation while ensuring the responsible advancement of LLM technologies.

### v3

@@ 5.4. Model: Deploy & Maintenance
Challenge I [SE Perspective]: Model Drift. The challenge of identifying and quantifying model drift in LLMs is considerable (Madaan et al., 2023a; Li et al., 2024c). LLMs are trained on vast datasets that are supposed to represent the linguistic diversity of their intended application domain. However, as the language evolves or the model is applied to slightly different contexts, the ability to remain relevant and consistent can shift in subtle ways. Recent research (Madaan et al., 2023a; Broscheit et al., 2022) emphasizes the need for sophisticated tools that can detect not only overt drifts in language usage but also more nuanced shifts in sentiment, context, or cultural references. These tools must be capable of parsing the complexities of human language, requiring ongoing refinement and adaptation to new linguistic phenomena.
  • Opportunity: Model Drift Monitoring. The realm of drift monitoring in LLMs presents a fertile ground for innovation and development. There is a significant opportunity to create and refine tools that can accurately detect and measure drift in various dimensions, from language usage to sentiment and contextual nuances. Furthermore, integrating these drift monitoring tools into the model development and deployment lifecycle can provide ongoing insights into model performance (Madaan et al., 2023a), enabling timely adjustments and enhancements. This proactive approach to managing model drift not only ensures the sustained relevance and accuracy of LLMs but also opens new avenues for research in understanding and mitigating the subtleties of language evolution in artificial intelligence.
@@ 6.1. Application: LLM Chatbots and Applications
Current State. Platforms like the GPT Store (OpenAI, 2024) are emerging as centralized hubs where developers can publish their LLM-powered apps (i.e., GPTs), and users can discover and access these tools to fulfill a wide array of tasks and objectives. This ecosystem draws parallels with mobile app stores (Fu et al., 2013; Lee and Raghu, 2014; Martin et al., 2017), aiming to create a curated, secure, and user-friendly environment for LLM-driven apps. The proliferation of LLM chatbots and apps is catalyzing innovation by lowering the barriers to entry for developers and providing them with platforms to build and deploy LLM-powered solutions. However, this rapid growth and accessibility also introduce new vulnerabilities and governance challenges. As these platforms evolve, they must grapple with novel security threats specific to LLM applications (Hou et al., 2024a; Su et al., 2024), such as prompt injection attacks and the potential misuse of advanced features like function calling. Furthermore, the unique nature of LLM apps, which can generate and manipulate content in real time, presents unprecedented challenges in terms of quality control, ethical considerations, and regulatory compliance.
Challenge I [S&P Perspective]: Security Vulnerabilities in LLM Apps. LLM-powered apps face significant security challenges, particularly in the realms of prompt injection attacks, misuse of function calling features, and prompt leakage (Wu et al., 2024a; OWASP, 2024b). Prompt injection attacks can manipulate the behavior of LLM chatbots, potentially leading to unauthorized actions or the disclosure of sensitive information (Liu et al., 2023a; Abdelnabi et al., 2023; Liu et al., 2023e). Function calling features, while powerful, introduce new attack vectors. Attackers may exploit these features to extract parameters and sensitive information about the underlying tools, hijack data flows, or conductresource exhaustion attacks (OWASP, 2024b) by overloading the system with malicious function calls (OWASP, 2024b). Additionally, prompt leakage poses a significant risk, where sensitive information embedded in prompts may be inadvertently revealed through the model’s responses, potentially compromising user privacy and system security (Hou et al., 2024a; Su et al., 2024).
@@ 6.2. Application: Autonomous LLM-based Agents
Current State. The current landscape of ALAs is characterized by rapid advancement and diverse applications across multiple domains. Recent developments have seen the emergence of more sophisticated agent architectures that combine LLMs with external tools, knowledge bases, and decision-making frameworks (Guo et al., 2024; Xi et al., 2023). These agents are increasingly capable of performing complex, multi-step tasks with minimal human intervention. For instance, in the field of software development, ALAs are being used for code generation, debugging, and even system design (Jin et al., 2024; Liu et al., 2024c; Wang et al., 2024d). However, alongside these advancements, there is growing concern about the ethical implications and potential risks associated with increasingly autonomous agents (He et al., 2024c; Li et al., 2024e; Zhang et al., 2024b).
Challenge I [SE Perspective]: Agent Optimization. From a software engineering perspective, a significant challenge in ALA development is the optimization of agent performance through efficient workflow design and orchestration (Sun et al., 2024b). This challenge falls under the emerging field of Software Engineering for LLM (SE4LLM). The complexity of tasks that ALAs can handle requires careful consideration of how different components interact, how information flows between stages of processing, and how to balance autonomy with controlled behavior. Effective agent optimization involves designing modular and reusable components, implementing efficient error handling and recovery mechanisms, and creating adaptive workflows that can adjust to varying task requirements and environmental conditions (Liu et al., 2024c; Wang et al., 2024d).
  • Opportunity: Workflow Orchestration Frameworks. This challenge presents an opportunity to develop sophisticated workflow orchestration frameworks specifically tailored for ALAs. Such frameworks could provide tools for visual workflow design, automated performance analysis, and dynamic workflow adjustment based on real-time feedback. Research in this area could focus on creating best practices for agent workflow design (Liu et al., 2024c), developing metrics for agent efficiency and effectiveness, and exploring ways to optimize the interaction between LLM components and external tools or data sources.
@@ 7. Conclusion
This paper presents the first comprehensive research agenda for the LLM supply chain, identifying the complex challenges and opportunities arising from the rapid advancement of large language models. By examining the LLM supply chain through the dual lenses of software engineering and security & privacy, we have provided a structured approach to understanding and advancing this critical field. Our analysis spans the infrastructure layer, foundation model layer, and downstream application ecosystem, presenting visions for robust and secure LLM development, reviewing current practices, and identifying key challenges and research opportunities. This work bridges the existing research gap in systematically understanding the multifaceted issues within the LLM supply chain. The proposed research agenda offers valuable insights to guide future efforts, fostering innovation while ensuring the responsible advancement of LLM technologies.
Acknowledgement
This work was supported in part by the Key R&D Program of Hubei Province (2023BAB017, 2023BAB079), the National Natural Science Foundation of China (grants No.62072046, 62302181), HUST CSE-HongXin Joint Institute for Cyber Security, HUST CSE-FiberHome Joint Institute for Cyber Security, and the Xiaomi Young Talents Program.